# Diffusion Implicit Policy for Unpaired Scene-aware Motion Synthesis

### v1

@@ 3.6 Multi-Task Motion Synthesis
4 Experiments
4.1 Datasets
Motion Datasets. Here, we use captured motion data from AMASS [24], which include action/description labels, to train our controlled motion diffusion model. Babel [31] provided action labels and the start/end frames for several subsets of AMASS. HumanML3D [12] provided additional sentence annotations and start/end frames for more motion data in AMASS. We match the keywords and categorize them into three states where details are presented in the Supplementary Material. All motions are downsampled to $40$ FPS and split into $160$-frame motion clips with a $20$-frame stride. Motion clips are all transformed according to the human pose in the first frame, where the transformed initial pose is centered (with the pelvis located at the origin) and orientation lie in $yz$-plane ($y\geq 0$). All the motion clips, along with the action labels, are used to train the motion diffusion model. Additionally, skeletons are extracted to provide joint position hints when training the ControlNet branch.
@@ 4.1 Datasets
SAMP [14] 5.97 0.14 0.84 0.94
GAMMA [57] 3.87 0.03 0.94 0.94
DIMOS [62] 6.43 0.04 0.99 0.95
@@
Table 1: Evaluation of motion synthesis on locomotion task. The up/down arrows ($\uparrow$/$\downarrow$) indicate higher/lower is better. Metrics with best performance are annotated in boldface.
Scene Datasets. We evaluate the performance of the proposed Diffusion Implicit Policy in both synthesized scenes and real scanned scenes. Following DIMOS [62], we use randomly generated scenes consisting of furniture from ShapeNet [6] to validate the performance on atomic locomotion and human-scene interaction. As for real scanned scenes from PROX [13] and Replica [38], we take them to evaluate the performance of the pipeline in synthesizing long-term motions within scenes that involve multiple tasks. All the experiments are conducted using the same controlled motion diffusion model and pipeline, thus indicating the generalization ability of the proposed method.
4.2 Scene Navigation
Figure 4: Visual results given by DIMOS and our method for locomotion task. The dashed circles indicate lower penetration, less skating and higher diversity in the synthesized motion.
Figure 5: Visual results of synthesized motions given by DIMOS and our method for sitting (left) and lying (right) task. The dashed circles indicates obvious advantages over DIMOS in less collision (col. 1,3), higher motion diversity (col. 2) and better foot contact (col. 4).
We take the generated scenes from DIMOS [62] for testing, where scenes are cluttered with furniture from ShapeNet [6]. In this experimental setting, the human need to walk from the starting point to the target point and avoid collisions with the furniture in scenes.
Metrics. We also evaluate the performance of synthesized motion for locomotion from four aspects as DIMOS [62], where (1) finish time measured in seconds, (2) average horizontal distance from final human body to target point measured in meters, (3) foot joint contact score
$s_{cont}=e^{-(|minj_{z}|-0.05)_{+}}\cdot e^{-(min||j_{vel}||_{2}-0.075)_{+}},$ (12)
and (4) locomotion penetration score indicating the percentage of body vertices within the walkable area are taken as our metrics.
Results. We compare our method with SAMP [14], GAMMA [57], and DIMOS [62] for locomotion in 3D scenes and report the results in Tab. 1. The results indicate our method can achieve the shortest finish time (3.35s), closest distance (0.03m), and lowest penetration (0.95). It’s noteworthy, the locomotion speed of the proposed method is much similar to that of real human than other methods. As can be seen, the performance on the contact score is inferior. We believe that is because our method focuses more on foot vertex contact, whereas the contact score calculation is based on foot joints.
@@ 4.2 Scene Navigation
4.3 Scene Object Interaction
We take scenes with furniture from ShapeNet [6] to evaluate the performance of the proposed method on scene object interaction, and $10$ objects ($3$ armchairs, $3$ straight chairs, $3$ sofas and $1$ L-sofa) are chosen for atomic interaction as previous work [62]. For each scene, the human is initialized to stand in front of the interaction object and guided to interact with it and finally return to original position.
Metrics. We take 4 metrics to evaluate the performance of interaction, including (1) the time to finish the task measured in seconds, (2) the foot contact score mentioned in Eq. 12, (3) mean human mesh vertex penetration $\sum_{v\in M}|(f_{SDF}(v))_{-}|$ over time, and (4) maximum penetration over time.
Results. We compare the proposed method with prevailing methods [14, 62] on human-scene interaction. The interaction tasks for sitting and lying are evaluate separately, and the results are reported in Tab. 2.
time $\downarrow$ contact $\uparrow$ pene. mean $\downarrow$ pene. max $\downarrow$
SAMP sit [14] 8.63 0.89 11.91 45.22
DIMOS sit [62] 4.09 0.97 1.91 10.61
Ours sit 3.71 0.89 1.86 7.13
SAMP lie [14] 12.55 0.73 44.77 238.81
DIMOS lie [62] 4.20 0.78 9.90 44.61
Ours lie 3.55 0.68 9.80 30.8
Table 2: Evaluation of motion synthesis on interaction tasks. The up/down arrows ($\uparrow$/$\downarrow$) indicate higher/lower is better. The best results are shown in boldface.
Fig. 5 visualizes the synthesized human-scene interaction given by DIMOS and the proposed method. The visual results indicate the proposed method present more plausible interaction with higher human-scene contact and lower human-scene collision.
4.4 Long-term Motion Synthesis
For long-term motion synthesis in 3D scenes where multiple tasks are completed consecutively, objects with feasible interaction in scenes are randomly selected. We utilize COINS [61] to sample the static interactions with these objects as the goals. All compared methods use the same initial state and task goals to synthesize long-term motions, ensuring a fair comparison.
@@ 4.4 Long-term Motion Synthesis
Results. Finally, $1,200$ ratings from $15$ participants are collected for each method ($60$ results per sample from $4$ aspects) and we report the scores in Tab. 3. It can be seen, the proposed method achieve the best performance in motion diversity, interaction plausibility and overall performance even with no need of paired motion-scene data for training. For motion naturalness, our proposed method is on par with DIMOS, as no specific designs are taken to further improve motion naturalness.
@@
Table 3: Comparison between competitive methods based on user study. Users are asked to give scores (ranging from $1$ to $5$, $\uparrow$) according to motion naturalness, diversity, interaction plausibility, and overall performance. Results on PROX/Replica are reported on the left/right respectively.
@@
We also show the results of scene-aware motion synthesis in scenes from PROX and Replica dataset respectively in Fig. 6 and Fig. 7. It can be seen that the proposed Diffusion Implicit Policy performs well in terms of motion naturalness, interaction plausibility and motion diversity thanks to the integration of motion denoising, implicit policy optimization, and random sampling within a unified framework.
Please kindly refer to the Supplementary Material for additional visual results, where ablation studies are also included.
5 Conclusion
@@
In this paper, we propose a unified framework, termed Diffusion Implicit Policy, for motion synthesis in 3D scenes, where paired motion-scene data is no longer needed for training. In this framework, interaction is disentangled from motion learning during training. Motion prior from diffusion model and implicit policy from reward functions are later integrated together to iteratively optimize the motion from random noise, pursuing motion naturalness, diversity and interaction plausibility simultaneously. We utilize joint hints and inpainting to ensure that keyframe poses remain consistent with the historical motion. We adjust the sample distribution centroid in a GAN Inversion manner to achieve better interaction plausibility while maintaining motion continuity. We introduce motion blending in the power space of the rotation matrix and the linear space of translation to ensure smooth transitions between multiple tasks for long-term motion synthesis. Comprehensive experiments on generated scenes with ShapeNet furniture, and scenes from PROX and Replica indicate the effectiveness and generalization capability. Such a promising solution can also encourages future works to learn scene-aware motion synthesis from unpaired motion and scene data.
Supplementary Material
6 Implementation Details
6.1 Motion Coordinate Transformation
@@
To reduce redundancy in motion representation for diffusion model training, we transform each motion into a local coordinate system.
@@
Specifically, given the original motion $x_{0}$ in world coordinates, we will fetch the pose in the first frame $P_{1}=\{\theta_{1,global},\theta_{1,j=1:21},\tau_{1}\}$. Next, we can obtain the corresponding human joints $J=J_{1:K}$ from the human pose $P_{1}$, human shape $\beta$, and hand pose $\theta_{h}$ using the SMPL-X human model [28]. In the SMPL-X human model, $J_{1:3}$ indicate the joints of the pelvis, left hip, and right hip.
@@
In the local coordinate system, the pelvis in the first frame should be positioned at the origin, thus we will first apply a translation $\mathbf{t}=-J_{1}$. Additionally, we need to apply a horizontal rotation $\mathbf{R}$ to ensure that the orientation of the human body in the first frame lie in the $yz$-plane ($y\geq 0$). When we only consider the local horizontal rotation for the initial human pose, we can set $\mathbf{z}_{l}=[0,0,1]$, $\mathbf{x}_{l}=\frac{\mathbf{n}}{||\mathbf{n}||}$ where $\mathbf{n}=[J_{3}[0]-J_{2}[0],J_{3}[1]-J_{2}[1],0]$ is the projection of $J_{3}-J_{2}$ on $xy$-plane, and $\mathbf{y}_{l}=\mathbf{z}_{l}\times\mathbf{x}_{l}$. At this point, the horizontal rotation can be formulated as
@@
Figure 8: More visual comparisons in 3D scenes from PROX.
Figure 9: Visual results of additional synthesized motions in 3D scenes from Replica dataset.
6.2 Reward Design
We take the following reward functions as an implicit policy for scene-aware motion synthesis to pursue motion continuity, goal achievement, and plausible human-scene interaction.
Historical Motion Consistency Reward. To ensure continuity with historical motion $\tilde{P}_{1:\tilde{S}}$ when inferring future motion $\hat{P}_{1:S}$, the poses in the first $H$ frames of $\hat{P}_{1:S}$ should be consistent with the last $H$ frames in historical motion $\tilde{P}_{\tilde{S}-H+1:\tilde{S}}$. Here, we impose constraints on the body joints to maintain consistency between historical and synthesized future motions. Such reward can be formulated as:
@@ 6.2 Reward Design
Smoothness Reward. During motion synthesis, we should also ensure the motion smoothness, and such reward is necessary when interaction-based implicit policy is taken into the diffusion model. Thus, we introduce a reward for limited acceleration in case of abrupt pose changes:
$\begin{array}[]{rcl}\mathcal{R}_{acc}&=&\sum_{s=2}^{S-1}(||M_{m,s+1}+M_{m,s-1}-2M_{m,s}||_{2}\\ &&\times\nu^{2}-\epsilon_{acc}).\end{array}$ (16)
@@
Goal Achievement Reward. We also encourage the synthesized motion to achieve the goal, thus we take the goal position or pose joints in the form of $J_{goal}$ as the guidance. Here, we decide the frame $g$ that need to be controlled according to the distance and sampled speed (which is correlated with the action state). The goal achievement reward for the synthesized motion is defined as:
@@
Contact Reward. The synthesized motion should keep in contact with the scene. For locomotion, at least one foot vertices should be in touch with the floor, thus the contact reward is define as:
$\begin{array}[]{rcl}\mathcal{R}_{cont}&=&\sum_{s=1}^{S}-ReLU(|\min_{f}(M_{f,s}[z])-h_{floor}|\\ &&-\epsilon_{cont}),\end{array}$ (18)
@@
$\begin{array}[]{rcl}\mathcal{R}_{cont}&=&\sum_{s=1}^{S}-ReLU(|\min_{m}(f_{SDF}(M_{m,s}))|\\ &&-\epsilon_{cont}).\end{array}$ (19)
Non-Penetration Reward. For the human-scene interaction in the synthesized motion, penetration should be avoided. We utilize the scene Signed Distance Function (SDF) as the guidance, and such reward function will encourage the generated human body move away from the interior space of the scene mesh. The non-penetration reward take the form of
@@
Here, $\epsilon_{pene}$ is the tolerance for slight penetration and we take SSM2 marker $M_{m}$ indicating $67$ body surface vertices to simplify the mesh like previous works [59, 62].
Non-Skating Reward. To avoid body skating during motion synthesis, we need to make sure that the velocity of contact body part to be close to 00. Here, we decide to fetch all possible contact parts (e.g., feet during walking, gluteus and back during lying). Thus, the reward for non-skating can be formulated as follows:
$\begin{array}[]{rcl}\mathcal{R}_{skt}&=&\sum_{s=1}^{S-1}-ReLU(\min_{c}(||M_{c,s+1}-M_{c,s}||_{2})\\ &&\times\nu-\epsilon_{vel}),\end{array}$ (21)
@@
As described in the main manuscript, the total reward function for implicit policy take the form
$\begin{array}[]{rcl}\mathcal{R}_{ip}&=&\lambda_{his}\mathcal{R}_{his}+\lambda_{goal}\mathcal{R}_{goal}+\lambda_{skt}\mathcal{R}_{skt}\\ &&+\lambda_{pene}\mathcal{R}_{pene}+\lambda_{cont}\mathcal{R}_{cont}\\ &&+\lambda_{acc}\mathcal{R}_{acc}.\end{array}$ (22)
Commonly, we set $\epsilon_{vel}=0.5$, $\epsilon_{pene}=0.03$, $\epsilon_{cont}=0.01$, and $\epsilon_{acc}=50$. For reward functions, $\lambda_{init}$ and $\lambda_{goal}$ are set to $1$, and $\lambda_{cont}$ is set to $10^{-1}$. For locomotion, $\lambda_{skt}$ is set to $10^{-3}$. For sitting, $\lambda_{skt}$, $\lambda_{pene}$, and $\lambda_{acc}$ are set to $3\times 10^{-4}$, $10^{-1}$, and $10^{-3}$. For lying, $\lambda_{pene}$ and $\lambda_{acc}$ are set to $3\times 10^{-2}$ and $10^{-3}$ respectively. All other coefficients are set to 00.
6.3 Motion Synthesis
The conditional motion diffusion model is designed to synthesize motion lasting $S=160$ frames within $T=10^{3}$ steps ($T_{inpaint}=50$). The first $K=22$ joints from SMPL-X model are selected to represent the body skeleton. Last $H$ (at most $H_{max}=10$) frames are taken as historical hints for motion blending. The keywords for action state categorization are shown in Tab. 4.
Action locomotion sit lie
Keywords walk, turn, jog, run sit lie, lying
7 Visualization Results
We present additional visual results on PROX and Replica in Fig. 8 and Fig. 9. Thanks to integration of motion denoising, implicit policy optimization, and random sampling within the proposed Diffusion Implicit Policy framework, motion naturalness, interaction plausibility and motion diversity can be obtained simultaneously.
8 Ablation Study
In this section, we conduct more experiments to validate the proposed framework and prove our claims.
8.1 Direct Optimization v.s. GAN Inversion
In the implicit policy optimization, we propose to optimize $\mu_{t}$ via $\hat{x}_{0}^{\varphi}(\mu_{t},t-1,c)$ rather than directly optimize $\mu_{t}$ as OmniControl [46]. In this way, a better sample distribution centroid $\mu_{t}$ can be searched in GAN Inversion manner to satisfy the interaction in final synthesized motion $\hat{x}_{0}^{\phi}$. In addition, $\mu_{t}$ can be optimized as a whole instead of optimizing only one pose in one frame, thus can better keep motion continuity. Here, we compare the synthesized motions of these two strategies to prove the advantage of the proposed framework on final motion naturalness. Fig. 10 illustrates the comparisons in a scene from PROX dataset, and we mark the synthesized motions with discontinuity in red/green circles. Without motion searching through GAN Inversion, the reward will directly guide the motion distribution centroid $\mu_{t}$, and we can see the sparse constrains in reward function will definitely make synthesized motions have abrupt changes.

### v2

@@ Introduction
Experiments
Table 1: Evaluation of motion synthesis on locomotion task. The up/down arrows ($\uparrow$$\downarrow$) indicate higher/lower is better. Metrics with best performance are annotated in boldface.
@@
SAMP 5.97 0.14 0.84 0.94
GAMMA 3.87 0.03 0.94 0.94
DIMOS 6.43 0.04 0.99 0.95
@@
Figure 4: Visual results given by DIMOS and ours for locomotion task. The dashed circles indicate lower penetration, less skating and higher diversity in the synthesized motion.
Figure 5: Visual results of synthesized motions given by DIMOS and our method for sitting and lying. The dashed circles indicates obvious advantages over DIMOS in less collision (col. 1,3), higher diversity (col. 2) and better foot contact (col. 4).
Datasets
Motion Datasets. Here, we use captured motion data with action/description labels from AMASS (mahmood2019amass) to train our model. Babel (punnakkal2021babel) provided action labels and the start/end frames for several subsets of AMASS. HumanML3D (guo2022generating) provided additional sentence annotations and start/end frames for more motion data in AMASS.
Scene Datasets. We evaluate the performance in both synthesized scenes and real scanned scenes. Following DIMOS (zhao2023synthesizing), we use randomly generated scenes consisting of furniture from ShapeNet (chang2015shapenet) to validate the performance on atomic locomotion and human-scene interaction. As for real scanned scenes from PROX (hassan2019resolving) and Replica (straub2019replica), we take them to evaluate the performance of the pipeline in synthesizing long-term motions within scenes that involve multiple tasks. All experiments are conducted using the same model and pipeline, thus indicating the generalization ability of our method.
Scene Navigation
We take the generated scenes from DIMOS (zhao2023synthesizing) for testing, where scenes are cluttered with furniture from ShapeNet (chang2015shapenet). In this experimental setting, the human need to walk from the starting point to the target point and avoid collisions with the furniture in scenes.
Metrics. We also evaluate the performance of synthesized motion for locomotion from four aspects as DIMOS (zhao2023synthesizing), where (1) finish time measured in seconds, (2) average horizontal distance from final human body to target point measured in meters, (3) foot joint contact score
[…]
and (4) penetration score indicating the percentage of body vertices within the walkable area are taken as our metrics.
Results. We compare our method with SAMP (hassan2021stochastic), GAMMA (zhang2022wanderings), and DIMOS (zhao2023synthesizing) for locomotion in 3D scenes and report the results in Tab. 1. The results indicate our method can achieve the shortest finish time (3.35s), closest distance (0.03m), and lowest penetration (0.95). It’s noteworthy, the locomotion speed of the proposed method is much similar to that of real human than other methods. As can be seen, the performance on the contact score is inferior. We believe that is because our method focuses more on foot vertex contact, whereas the contact score calculation is based on foot joints.
@@
Scene Object Interaction
We take scenes with furniture from ShapeNet (chang2015shapenet) to evaluate the performance of the proposed method on scene object interaction, and $10$ objects ($3$ armchairs, $3$ straight chairs, $3$ sofas and $1$ L-sofa) are chosen for atomic interaction as previous work (zhao2023synthesizing). For each scene, the human is initialized to stand in front of the interaction object and guided to interact with it and finally return to original position.
Metrics. We take 4 metrics to evaluate the performance, including (1) the time to finish the task measured in seconds, (2) the foot contact score mentioned in Eq. 12, (3) mean human mesh vertex penetration $\sum_{v\in M}|(f_{SDF}(v))_{-}|$ over time, and (4) maximum penetration over time.
Results. We compare the proposed method with prevailing methods (hassan2021stochastic; zhao2023synthesizing) on human-scene interaction. The interaction tasks for sitting and lying are evaluate separately, and the results are reported in Tab. 2.
Table 2: Evaluation of motion synthesis on interaction tasks. The up/down arrows ($\uparrow$$\downarrow$) indicate higher/lower is better. The best results are shown in boldface.
time $\downarrow$ pene. mean $\downarrow$ pene. max $\downarrow$
SAMP sit 8.63 11.91 45.22
DIMOS sit 4.09 1.91 10.61
Ours sit 3.71 1.86 7.13
SAMP lie 12.55 44.77 238.81
DIMOS lie 4.20 9.90 44.61
Ours lie 3.55 9.80 30.8
Fig. 5 visualizes the synthesized human-scene interaction given by DIMOS and our method. The visual results indicate our method presents more plausible interaction with higher human-scene contact and lower human-scene collision.
@@
Long-term Motion Synthesis
For long-term motion synthesis in 3D scenes where multiple tasks are completed consecutively, objects with feasible interaction in scenes are randomly selected. We utilize COINS (zhao2022compositional) to sample the static interactions with these objects as the goals. All compared methods use the same initial state and task goals to synthesize long-term motions, ensuring a fair comparison.
@@
Results. Finally, $1,200$ ratings from $15$ participants are collected for each method ($60$ results per sample from $4$ aspects) and we report the scores in Tab. 3. It can be seen, the proposed method achieve the best performance in motion diversity, interaction plausibility and overall performance even without paired motion-scene data for training. For motion naturalness, our proposed method is on par with DIMOS, as no specific designs are taken to improve motion naturalness.
Table 3: Comparison between competitive methods based on user study. Users are asked to give scores (ranging from 11 to $5$, $\uparrow$) according to motion naturalness, diversity, interaction plausibility, and overall performance. Results on PROX/Replica are reported on the left/right respectively.
@@
We present the results of motion synthesis in scenes from PROX and Replica dataset respectively in Fig. 6 and Fig. 7. It can be seen that the proposed Diffusion Implicit Policy performs better in terms of motion naturalness, interaction plausibility and motion diversity. This is due to the integration of motion denoising, implicit policy optimization, and random sampling within a unified framework.
Conclusion
@@
In this paper, we propose a unified framework, termed Diffusion Implicit Policy, for motion synthesis in 3D scenes, where paired motion-scene data is no longer needed for training. In this framework, interaction is disentangled from motion learning during training. Motion prior from diffusion model and implicit policy from reward functions are integrated together to iteratively optimize the motion from random noise, pursuing motion naturalness, diversity and interaction plausibility simultaneously. We utilize joint hints and inpainting to ensure that keyframe poses remain consistent with the historical motion. We adjust the motion distribution centroid in a GAN Inversion manner to achieve better interaction plausibility while maintaining motion continuity. We introduce motion blending in the power space of the joint rotation matrix to ensure smooth transitions between multiple tasks for long-term motion synthesis. Comprehensive experiments on generated scenes with ShapeNet furniture, and scenes from PROX and Replica indicate the effectiveness and generalization capability. Such a promising solution can also encourages future works to learn scene-aware motion synthesis from unpaired motion and scene data.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (Grant No. 62222602, 62302167, U23A20343, 62502159), Natural Science Foundation of Shanghai (Grant No. 25ZR1402135), Shanghai Sailing Program (Grant No. 23YF1410500), Young Elite Scientists Sponsorship Program by CAST (Grant No. YESS20240780), the Chenguang Program of Shanghai Education Development Foundation and Shanghai Municipal Education Commission (Grant No. 23CGA34), Natural Science Foundation of Chongqing (Grant No. CSTB2023NSCQ-JQX0007,CSTB2023NSCQ-MSX0137, CSTB2025NSCQ-GPX0445), Open Project Program of the State Key Laboratory of CAD&CG (Grant No. A2501), Zhejiang University.
Appendix A Implementation Details
In this section, we will provide implementation details on motion data preparation, motion coordinate transformation, reward design and motion synthesis procedure.
Data Preparation
We utilize the motion data from AMASS (mahmood2019amass) for training. Specifically, Babel (punnakkal2021babel) provided action labels and HumanML3D (guo2022generating) provided additional sentence annotations for several subsets of the AMASS. For these motions, we match the keywords and categorize them into three states. The keywords for action state categorization are shown in Tab. 4. All motions are downsampled to $40$ FPS and split into $160$-frame motion clips with a $20$-frame stride. Motion clips are all transformed according to the human pose in the first frame, where the transformed initial pose is centered (with the pelvis located at the origin) and orientation lie in $yz$-plane ($y\geq 0$). All the motion clips, along with the action labels, are used to train the motion diffusion model. Additionally, skeletons are extracted to provide joint position hints when training the ControlNet branch.
Action locomotion sit lie
Keywords walk, turn, jog, run sit lie, lying
Motion Coordinate Transformation
@@ Appendix A Implementation Details
To reduce redundancy in motion representation for diffusion model training, we transform each motion clip into a local coordinate system.
@@
Specifically, given the original motion $x_{0}$ in world coordinates, we will fetch the pose in the first frame $P_{1}=\{\theta_{1,global},\theta_{1,j=1:21},\tau_{1}\}$. Next, we can obtain the corresponding human joints $J=J_{1:K}$ from the human pose $P_{1}$, human shape $\beta$, and hand pose $\theta_{h}$ using the SMPL-X human model (pavlakos2019expressive). In the SMPL-X human model, $J_{1:3}$ indicate the joints of the pelvis, left hip, and right hip.
@@
In the local coordinate system, the pelvis in the first frame should be positioned at the origin, thus we will first apply a translation $\mathbf{t}=-J_{1}$. Additionally, we need to apply a horizontal rotation $\mathbf{R}$ to ensure that the orientation of the human body in the first frame lie in the $yz$-plane ($y\geq 0$). When we only consider the local horizontal rotation for the initial human pose, we can set $\mathbf{z}_{l}=[0,0,1]$, $𝐧\mathbf{x}_{l}$$\frac$$\mathbf$$|\mathbf{n}$||} where $\mathbf{n}=[J_{3}[0]-J_{2}[0],J_{3}[1]-J_{2}[1],0]$ is the projection of $J_{3}-J_{2}$ on $xy$-plane, and $\mathbf{y}_{l}=\mathbf{z}_{l}\times\mathbf{x}_{l}$. At this point, the horizontal rotation can be formulated as
@@
Reward Design
In our implementation, Eq. 5 in the main manuscript take the following reward functions as an implicit policy for scene-aware motion synthesis to pursue motion continuity, goal achievement, and plausible human-scene interaction.
Historical Motion Consistency Reward.
To ensure continuity with historical motion $\tilde{P}_{1:\tilde{S}}$ when inferring future motion $\hat{P}_{1:S}$, the poses in the first $H$ frames of $\hat{P}_{1:S}$ should be consistent with the last $H$ frames in historical motion $\tilde{P}_{\tilde{S}-H+1:\tilde{S}}$. Here, we impose constraints on the body joints to maintain consistency between historical and synthesized future motions. Such reward can be formulated as:
@@
Smoothness Reward.
During motion synthesis, we should also ensure the motion smoothness, and such reward is necessary when interaction-based implicit policy is taken into the diffusion model. Thus, we introduce a reward for limited acceleration in case of abrupt pose changes:
$\begin{array}[]{rcl}\hskip-2.84544pt\mathcal{R}_{acc}&\hskip-8.5359pt=&\hskip-8.5359pt\sum_{s=2}^{S-1}(||M_{m,s+1}+M_{m,s-1}-2M_{m,s}||_{2}\\ &&\hskip-8.5359pt\times\nu^{2}-\epsilon_{acc}).\end{array}$ (16)
@@
Goal Achievement Reward.
We also encourage the synthesized motion to achieve the goal, thus we take the goal position or pose joints in the form of $J_{goal}$ as the guidance. Here, we decide the frame $g$ that need to be controlled according to the distance and sampled speed (which is correlated with the action state). The goal achievement reward for the synthesized motion is defined as:
@@
Contact Reward.
The synthesized motion should keep in contact with the scene. For locomotion, at least one foot vertices should be in touch with the floor, thus the contact reward is define as:
$\begin{array}[]{rcl}\hskip-2.84544pt\mathcal{R}_{cont}&\hskip-8.5359pt=&\hskip-8.5359pt\sum_{s=1}^{S}-ReLU(|\min_{f}(M_{f,s}[z])-h_{floor}|\\ &&\hskip-8.5359pt-\epsilon_{cont}),\end{array}$ (18)
@@
$\begin{array}[]{rcl}\hskip-2.84544pt\mathcal{R}_{cont}&\hskip-8.5359pt=&\hskip-8.5359pt\sum_{s=1}^{S}-ReLU(|\min_{m}(f_{SDF}(M_{m,s}))|\\ &&\hskip-8.5359pt-\epsilon_{cont}).\end{array}$ (19)
Non-Penetration Reward.
For the human-scene interaction in the synthesized motion, penetration should be avoided. We utilize the scene Signed Distance Function (SDF) as the guidance, and such reward function will encourage the generated human body move away from the interior space of the scene mesh. The non-penetration reward take the form of
@@
Here, $\epsilon_{pene}$ is the tolerance for slight penetration and we take SSM2 marker $M_{m}$ indicating $67$ body surface vertices to simplify the mesh like previous works (zhang2021we; zhao2023synthesizing).
Non-Skating Reward.
To avoid body skating during motion synthesis, we need to make sure that the velocity of contact body part to be close to 0. Here, we decide to fetch all possible contact parts (e.g. , feet during walking, gluteus and back during lying). Thus, the reward for non-skating can be formulated as follows:
$\begin{array}[]{rcl}\mathcal{R}_{skt}&\hskip-8.5359pt=&\hskip-8.5359pt\sum_{s=1}^{S-1}-ReLU(\min_{c}(||M_{c,s+1}-M_{c,s}||_{2})\\ &&\hskip-8.5359pt\times\nu-\epsilon_{vel}),\end{array}$ (21)
@@
Commonly, we set $\epsilon_{vel}=0.5$, $\epsilon_{pene}=0.03$, $\epsilon_{cont}=0.01$, and $\epsilon_{acc}=50$. For reward functions, $\lambda_{init}$ and $\lambda_{goal}$ are set to $1$, and $\lambda_{cont}$ is set to $10^{-1}$. For locomotion, $\lambda_{skt}$ is set to $10^{-3}$. For sitting, $\lambda_{skt}$, $\lambda_{pene}$, and $\lambda_{acc}$ are set to $3\times 10^{-4}$, $10^{-1}$, and $10^{-3}$. For lying, $\lambda_{pene}$ and $\lambda_{acc}$ are set to $3\times 10^{-2}$ and $10^{-3}$ respectively. All other coefficients are set to 0.
Motion Synthesis
The conditional motion diffusion model is designed to synthesize motion lasting $S=160$ frames within $T=10^{3}$ steps ($T_{inpaint}=50$). The first $K=22$ joints from SMPL-X model are selected to represent the body skeleton. Last $H$ (at most $H_{max}=10$) frames are taken as historical hints for motion blending.
Computing Infrastructure
All experiments can be conducted on a RTX 3090 with 64G RAM. We also provide the explicit versions of software libraries in the supplementary code.
Appendix B More Visualization Results
In this section, we present more visualization results of our method in scenes from PROX (hassan2019resolving) and Replica (straub2019replica). In Fig. 8, we provide additional visual comparison of different methods in PROX dataset. In Fig. 9, we visualize more results of competitive methods on Replica dataset.
Figure 8: Additional visual comparison of different methods on motion synthesis in 3D scenes from PROX.
Figure 9: More visual results of synthesized motions given by compared methods in Replica scenes.
Ablation Study
Here, we conduct more experiments to validate the proposed framework and prove our claims.
Direct Optimization v.s. GAN Inversion. In the implicit policy optimization, we propose to optimize $\mu_{t}$ via $\hat{x}_{0}^{\varphi}(\mu_{t},t-1,c)$ rather than directly optimize $\mu_{t}$ as OmniControl (xie2024omnicontrol). In this way, a better sample distribution centroid $\mu_{t}$ can be searched in GAN Inversion manner to satisfy the interaction in final synthesized motion $\hat{x}_{0}^{\phi}$. In addition, $\mu_{t}$ can be optimized as a whole instead of optimizing only one pose in one frame, thus can better keep motion continuity. Here, we compare the synthesized motions of these two strategies to prove the advantage of the proposed framework on final motion naturalness. Fig. 10 illustrates the comparisons in a scene from PROX dataset, and we mark the synthesized motions with discontinuity in red/green circles. Without motion searching through GAN Inversion, the reward will directly guide the motion distribution centroid $\mu_{t}$, and we can see the sparse constrains in reward function will definitely make synthesized motions have abrupt changes.